# A matrix-free convex limiting framework for continuous Galerkin methods with nonlinear stabilization

### v1

@@ 5.1.1 Linear advection
To begin, we advect the $L^{2}(\Omega)$ projection $u_{h}(\cdot,0)\in V_{h}$ of the smooth initial profile $u_{0}(x)=\mathrm{e}^{-100(x-0.5)^{2}}$ to study the convergence behavior of the CG-WENO target scheme (10) and of its limited counterpart (38). The $L^{2}(\Omega)$ errors at the final time $t=1$ and the corresponding estimated orders of convergence (EOC) for $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$ elements are listed in Table˜1.
@@
While the optimal convergence rate $p+1$ is attained or even exceeded by the target scheme for Bernstein finite elements of degree $p\in\{1,2\}$, it turns out that even slope limiting w. r. t. global bounds is too restrictive in the $\mathbb{P}_{2}$ case. The issue here is subtle and can be explained as follows: While the pointwise values of the exact solution remain within the range $\mathcal{G}=[0,1]$, this is not necessarily the case for the Bernstein coefficients. This peculiarity of Bernstein-basis AFC schemes is the reason why constraining the degrees of freedom to stay in $\mathcal{G}$ leads to peak clipping effects and a loss of accuracy in the neighborhood of smooth global extrema (cf. [17]). To achieve optimal convergence, we envision the use of collocated LGL spectral elements (cf. [47, 51]) in future implementations of the proposed methodology.
@@
We observed similar convergence behavior in numerical experiments for the nonlinear Burgers test from [36, Sec. 7.2]. The results (not shown here) confirm that the target scheme converges at the optimal rate $p+1$. The limited $\mathbb{P}_{1}$ approximation ($p=1$) exhibits second-order convergence, but the imposition of global bounds on the Bernstein coefficients of the $\mathbb{P}_{2}$ version prevented us from achieving full third-order accuracy.
In the next linear advection test, we advect the initial profile [14]
@@
The results shown in Figs.˜1(a) and 1(b) confirm that flux limiting can drastically mitigate the time step restrictions. Interestingly enough, the flux-corrected approximations even seem to be less diffusive than numerical solutions obtained without flux limiting but using smaller time steps. This behavior indicates that smearing effects tend to become less pronounced if fewer time steps are needed to reach the final time. It can also be observed that the low-order results become less accurate as the polynomial degree is increased while the mesh is coarsened to match the fixed value of $N_{h}$. In view of the findings reported in [16, 34, 40], this behavior is caused by the fact that the inherent Rusanov dissipation (15) lacks sparsity.
In our experience, the accuracy of the low-order solutions can be greatly improved by using the largest fake time step $\Delta t_{e}$ that ensures the IDP property of $\bar{u}^{e}$ for the given data. However, the results presented in Fig.˜1(a) show that the accuracy of slope-limited high-order approximations is not significantly affected by the levels of numerical dissipation in the low-order method corresponding to $\beta_{e}=0$. As expected, the step function is slightly better captured with $\mathbb{P}_{1}$ elements, while the $\mathbb{P}_{2}$ version preserves the peak of the smooth bump with higher precision, despite the fact that the corresponding mesh is one level coarser than that for the $\mathbb{P}_{1}$ simulation.
While it is remarkable that flux limiting ensures the validity of IDP constraints under the large-cell CFL condition, the requirement of stability for the unconstrained target (CG-WENO) scheme may dictate the use of smaller time steps in practice. If the target is unstable, violations of global bounds can be prevented by flux/slope limiters, but spurious oscillations may occur within these bounds, leading to unbounded growth of total variation or approximations of poor quality (see, e. g., [15, left panel of Fig. 3]). The CFL bound of the high-order scheme using Bernstein finite elements does depend on the polynomial degree and may be just marginally larger than the time step that meets the requirements of Lemma 1. Moreover, the efficiency gain resulting from the use of slightly larger $\Delta t$ is diminished by the overhead cost associated with flux limiting. For spectral element approximations of LGL type, the CFL condition of the high-order discretization is far less severe than the IDP time step restriction (27) of Lemma 1. Hence, the possibility of optional flux limiting may lead to substantial speedups in the LGL version, which represents a promising alternative to Bernstein polynomials.
@@ 5.1.2 Nonlinear problem with a nonconvex flux
Next, we consider the first of the two Riemann problems proposed in [23, Sec. 3.1] for the 1D nonlinear conservation law with the $C^{1}(\mathbb{R})$ flux function
@@
The domain $\Omega=(0,1)$ has an inlet at $x=0$ and an outlet at $x=1$. The inflow boundary condition $u(0,t)=0$ is prescribed in this test. The initial condition is given by $u_{0}(x)=\chi_{\{x\geq 1/4\}}$, where $\chi_{X}$ is the characteristic function of a set $X\subset\Omega$. The unique entropy solution consists of a shock propagating to the right and a rarefaction wave that connects the pre-shock value with the right state of the Riemann problem. A closed-form expression for the vanishing viscosity solution can be found in [27, Suppl. materials, Sec. A.4.2.2]. Many numerical methods fail to predict the pre-shock value correctly and yield profiles with piecewise-constant areas around the shock.
As in the previous example, we first use the low order methods to solve this problem up to the final time $t=1$ using a total of $N_{h}=129$ nodes for all combinations of polynomial degree $p$ and mesh resolution $h$. The low-order results obtained without and with flux limiting are displayed in Figs.˜2(a) and 2(b). The slope-limited recovery of auxiliary nodal states $\bar{u}_{i}^{e}$ from the unlimited high-order version of $\bar{u}^{e}$ yields the approximations shown in Fig.˜2(c) for $p\in\{1,2\}$. Since the exact solution is piecewise linear in this test, the $\mathbb{P}_{1}$ solution is more accurate than the $\mathbb{P}_{2}$ approximation with the same number of nodes. However, the accuracy of the latter is still satisfactory.
@@ 5.2 Euler equations in 1D
Let us now consider the 1D compressible Euler equations, a nonlinear hyperbolic system of conservation laws for the density $\rho$, momentum $\rho v$, and total energy $\rho E$ of an ideal gas. The pressure $p$ is given by the equation of state $p=(\gamma-1)(\rho E-\tfrac{1}{2}\rho|v|^{2})$, where $\gamma=1.4$ is the adiabatic constant. An invariant domain $\mathcal{G}$ of the Euler system consists of all states $u=(\rho,\rho v,\rho E)^{\top}$ such that $\rho>0$ and $p(u)>0$. The linear positivity constraint for $\rho$ can be enforced as in the scalar case. To ensure positivity preservation for $p(u)$, we adapt the limiter proposed by Abgrall et al. [3, 57] to our purposes.
Our numerical tests for the Euler equations are restricted to 1D in this work. We intend to perform 2D tests and extend the proposed methodology to high-order LGL spectral elements in a subsequent publication.
@@ 5.2.1 Modified Sod’s shock tube
In the first test for the Euler equations, we consider a 1D Riemann problem corresponding to a modification [55, Sec. 6.4] of Sod’s shock tube problem [54]. The modified version is more challenging because a sonic point occurs within the rarefaction wave region. As a consequence, numerical methods that lack entropy stability tend to produce entropy shocks.
@@
The left boundary of the domain $\Omega=(0,1)$ is a supersonic inlet, at which the values of the primitive variables are determined by the left initial state of the Riemann problem. The boundary data $\hat{u}$ for $x=1$ may be set equal to the right initial state, since the final time $t=0.2$ is too short for the waves to reach the right boundary. The exact entropy solution features a rarefaction, a contact, and a shock wave. Its derivation can be found in  [55, Ch. 4].
We solve this test problem with the target scheme and its slope-limited counterpart using $N_{h}=129$ degrees of freedom per variable and $p\in\{1,2\}$. The results displayed in Figs.˜3(a) and 3(b) are in good agreement with the vanishing viscosity solution. Similarly to the test problem considered in Section˜5.1.2, the use of $\mathbb{P}_{1}$ elements and fine grids is the best option in this example. While higher-order approximations on coarser grids perform well for rarefaction waves, they are typically more diffusive at discontinuities.
@@
In Fig.˜3(c), we plot the evolution of $\eta_{\Omega}(u)=\frac{1}{1-\gamma}\int_{\Omega}\rho s(u)\mathrm{d}x$, where $s(u)=\log\left(p\rho^{-\gamma}\right)$ is the specific physical entropy. In contrast to the test case considered in Section˜5.1.2, the entropy flux $q(u)=\rho vs/(1-\gamma)$ vanishes at the boundaries, as long as the states $u(0,t)$ and $u(1,t)$ remain equal to the corresponding initial states of the Riemann problem. On the time interval for which this is the case, the total entropy $\eta_{\Omega}(u)$ of the exact vanishing viscosity solution $u$ is subject to change only due to entropy losses at the shock and/or contact. Numerically, we observe a general trend of entropy dissipation but $\eta_{\Omega}(u_{h})$ does not decrease monotonically. A probable cause of this behavior is the lack of local entropy stability, which can be cured by adding entropy viscosity or using limiter-based entropy fixes [32]. However, the small wiggles in the evolution of $\eta_{\Omega}(u_{h})$ are no major concern, since no entropy shocks are observed and $\eta_{\Omega}(u_{h})\leq\eta_{\Omega}(u)$. As demonstrated already in [36] in the context of scalar problems, the WENO stabilization term of our method dissipates enough entropy in most cases. Although additional entropy stabilization may still be appropriate for some scalar problems (such as the challenging but artificial KPP test in Section 5.3.2), we have never observed entropy-violating behavior of IDP finite element schemes for the Euler equations.
@@ 5.2.2 Woodward–Colella blast wave problem
The 1D blast wave problem [58] is a challenging test for the Euler equations due to strong jumps in the piecewise-constant initial pressure
@@
The initial density is $\rho_{0}\equiv 1$, and the fluid is at rest, i. e., $v_{0}\equiv 0$. The domain $\Omega=(0,1)$ has reflecting boundaries at both end points. For short time intervals, the exact solution can be obtained as for any Riemann problem of the Euler equations [55, Ch. 4]. However, by the final time $t=0.038$, interactions with the boundaries and collisions of waves generated by the two Riemann problems result in the unavailability of an exact solution.
We solve this test problem numerically using $N_{h}=1001$ degrees of freedom per variable and $p\in\{1,2\}$. The density profile calculated with the target scheme is shown in Fig.˜4(a). Since no slope limiting is performed in this experiment, some negative pressures arise during the simulation run. The code does not crash only because we calculate the speed of sound using the formula $c=\sqrt{\max\{0,\gamma p/\rho\}}$. This practice is not to be recommended for non-IDP schemes, since it gives rise to consistency errors and disguises unacceptable pressure behavior. Due to round-off errors, a numerical implementation of an IDP scheme might also produce negative pressures of the order of machine precision. In this case, setting $c\coloneqq 0$ is an acceptable way to avoid complex-valued solutions. However, such situations did not occur in simulations in which we used the density limiter and the pressure limiter proposed by Abgrall et al. [3]. The slope-limited density profiles shown in Fig.˜4(b) are nonoscillatory, and no negative pressures were generated at any time.
@@ 5.2.3 Shu–Osher test / sine-shock interaction
Next, we perform simulations for the popular Shu–Osher test problem [53], which uses the initial condition
@@
We test the methods under investigation using $p\in\{1,2\}$ and $N_{h}=501$ degrees of freedom per variable. Since the exact solution can be captured well by the target scheme, no violations of IDP constraints occur during the simulation run without the slope limiter. Therefore, the density and pressure fixes are not activated in practice, and the results coincide with those obtained without limiting (see Fig.˜5).
It is quite encouraging to see that the $\mathbb{P}_{2}$ version captures the post-shock oscillatory region much better than the $\mathbb{P}_{1}$ approximation on a mesh with twice as many cells. Furthermore, as is the case for the 1D blast wave problem studied in Section˜5.2.2, limiting w. r. t. global bounds does not lead to severe peak clipping effects. However, one can easily think of situations in which this issue is likely to occur, e. g., in simulations of near-vacuum states with quadratic ($\mathbb{P}_{2}$) or higher order Bernstein elements. As long as the quantities to be limited remain sufficiently far away from the global bounds of IDP constraints, optimal convergence can be achieved with high-order Bernstein elements. In the context of AFC schemes for the advection equation, numerical evidence for the validity of this claim was provided in [17, Sec. 7.1].
@@ 5.3 Scalar problems in 2D
To evaluate the numerical behavior of the limited CG-WENO scheme and its components in two space dimensions, we consider two additional scalar test problems, one of which is linear and the other one is nonlinear. In both cases, we restrict our studies to $\mathbb{Q}_{1}$ finite element approximations, for which the Bernstein and Lagrange basis functions coincide. Computations are performed on uniform meshes with spacing $h=\frac{1}{128}$. We use the fake time step $\Delta t_{e}=\Delta t_{e}^{\max}$ defined by (37) and the global time step $\Delta t=10^{-3}$, which satisfies the CFL condition (23) for both test problems. Although our choice of $\Delta t_{e}$ exceeds the threshold of Lemma 1, the high-order intermediate cell averages stay in bounds. Therefore, we deactivate the flux limiter for (41) by setting $\alpha_{ee^{\prime}}\coloneqq 1\ \forall e\in\mathcal{B}_{e}$, as proposed in Remark 1.
@@
where $\dot{u}_{h}^{L}$ denotes a low-order approximation to the time derivative $\dot{u}_{h}$. This choice is adopted because it corresponds to the second-order target employed in [25]. A decrease in the value of $\gamma_{e}\in[0,1]$ on the right-hand side of (55) increases the levels of low-order Rusanov dissipation (15), while decreasing the levels of high-order stabilization via the terms depending on $\Delta t_{e}$.
@@
In the WENO version, we use the linear weights $w_{l}^{e,\rm lin}=0.2$ for $l=1,\ldots,,n_{e}$ and set $w_{0}^{e,\rm lin}=1-\sum_{l=1}^{n_{e}}w_{l}^{e,\rm lin}$. To study the dependence of the results on the parameter $q\geq 1$ of formula (7), we run simulations for $q\in\{3,5,10\}$. The range of approximate solutions is reported in the figures to quantify the levels of numerical diffusion and the magnitude of undershoots/overshoots.

### v2

@@ 5.1.1 Linear advection
To begin, we advect the $L^{2}(\Omega)$ projection $u_{h}(\cdot,0)\in V_{h}$ of the smooth initial profile $u_{0}(x)=\mathrm{e}^{-100(x-0.5)^{2}}$ to study the convergence behavior of the CG-WENO target scheme (10) and of its limited counterpart (41). The $L^{2}(\Omega)$ errors at the final time $t=1$ and the corresponding experimental orders of convergence (EOC) for $\mathbb{P}_{1}$ and $\mathbb{P}_{2}$ elements are listed in Table˜2.
@@
While the optimal convergence rate $p+1$ is attained or even exceeded by the target scheme for Bernstein finite elements of degree $p\in\{1,2\}$, it turns out that even slope limiting w. r. t. global bounds is too restrictive in the $\mathbb{P}_{2}$ case. The issue here is subtle and can be explained as follows: While the pointwise values of the exact solution remain within the range $\mathcal{G}=[0,1]$, this is not necessarily the case for the Bernstein coefficients. This peculiarity of Bernstein-basis AFC schemes is the reason why constraining the degrees of freedom to stay in $\mathcal{G}$ leads to peak clipping effects and a loss of accuracy in the neighborhood of smooth global extrema (cf. [hajduk2020b]). To achieve optimal convergence, we envision the use of collocated LGL spectral elements (cf. [pazner2021, rueda-ramirez2024]) in future implementations of the proposed methodology.
@@
We observed similar convergence behavior in numerical experiments for the nonlinear Burgers test from [kuzmin2023a, Sec. 7.2]. The results (not shown here) confirm that the target scheme converges at the optimal rate $p+1$. The limited $\mathbb{P}_{1}$ approximation exhibits second-order convergence, but the imposition of global bounds on the Bernstein coefficients of the $\mathbb{P}_{2}$ version prevented us from achieving full third-order accuracy.
In the next linear advection test, we advect the initial profile [hajduk2021]
@@
The results shown in Figs.˜2(a) and 2(b) confirm that flux limiting can drastically mitigate the time step restrictions. Interestingly enough, the flux-corrected approximations even seem to be less diffusive than numerical solutions obtained without flux limiting but using smaller time steps. This behavior indicates that smearing effects tend to become less pronounced if fewer time steps are needed to reach the final time. It can also be observed that the low-order results become less accurate as the polynomial degree is increased while the mesh is coarsened to match the fixed value of $N_{h}$. In view of the findings reported in [hajduk2020, kuzmin2020a, lohmann2017], this behavior is caused by the fact that the inherent Rusanov dissipation (15) lacks sparsity.
In our experience, the accuracy of the low-order solutions can be greatly improved by using the largest pseudo time step $\Delta t_{e}$ that ensures the IDP property of $\bar{u}^{e}$ for the given data. However, the results presented in Fig.˜2(a) show that the accuracy of slope-limited high-order approximations is not significantly affected by the levels of numerical dissipation in the low-order method corresponding to $\beta_{e}=0$. As expected, the step function is slightly better captured with $\mathbb{P}_{1}$ elements, while the $\mathbb{P}_{2}$ version preserves the peak of the smooth bump with higher precision, despite the fact that the corresponding mesh is one level coarser than that for the $\mathbb{P}_{1}$ simulation.
While it is remarkable that flux limiting ensures the validity of IDP constraints under the large-cell CFL condition, the requirement of stability for the unconstrained target (CG-WENO) scheme may dictate the use of smaller time steps in practice. If the target is unstable, violations of global bounds can be prevented by flux/slope limiters, but spurious oscillations may occur within these bounds, leading to unbounded growth of total variation or approximations of poor quality (see, e. g., [hajduk2025a, left panel of Fig. 3]). The CFL bound of the high-order scheme using Bernstein finite elements does depend on the polynomial degree and may be just marginally larger than the time step that meets the requirements of Lemma 2. Moreover, the efficiency gain resulting from the use of slightly larger $\Delta t$ is diminished by the overhead cost associated with flux limiting. For spectral element approximations of LGL type, the CFL condition of the high-order discretization is far less severe than the IDP time step restriction (29) of Lemma 2. Hence, the possibility of optional flux limiting may lead to substantial speedups in the LGL version, which represents a promising alternative to Bernstein polynomials.
@@ 5.1.2 Nonlinear problem with a nonconvex flux
Next, we consider the first of the two Riemann problems proposed in [kurganov2007, Sec. 3.1] for the 1D nonlinear conservation law with the $C^{1}(\mathbb{R})$ flux function
@@
The domain $\Omega=(0,1)$ has an inlet at $x=0$ and an outlet at $x=1$. The inflow boundary condition $u(0,t)=0$ is prescribed in this test. The initial condition is given by $u_{0}(x)=\chi_{\{x\geq 1/4\}}$, where $\chi_{X}$ is the characteristic function of a set $X\subset\Omega$. The unique entropy solution consists of a shock propagating to the right and a rarefaction wave that connects the pre-shock value with the right state of the Riemann problem. A closed-form expression for the vanishing viscosity solution can be found in [kuzmin2023, Suppl. materials, Sec. A.4.2.2]. Many numerical methods fail to predict the pre-shock value correctly and yield profiles with piecewise-constant areas around the shock.
As in the previous example, we first use the low-order methods to solve this problem up to the final time $t=1$ using a total of $N_{h}=129$ nodes for all combinations of polynomial degree $p$ and mesh resolution $h$. The low-order results obtained without and with flux limiting are displayed in Figs.˜3(a) and 3(b). The slope-limited recovery of auxiliary nodal states $\bar{u}_{i}^{e}$ from the unlimited high-order version of $\bar{u}^{e}$ yields the approximations shown in Fig.˜3(c) for $p\in\{1,2\}$. Since the exact solution is piecewise linear in this test, the $\mathbb{P}_{1}$ solution is more accurate than the $\mathbb{P}_{2}$ approximation with the same number of nodes. However, the accuracy of the latter is still satisfactory.
@@ 5.2 Euler equations in 1D
Let us now consider the 1D compressible Euler equations, a nonlinear hyperbolic system of conservation laws for the density $\rho$, momentum $\rho v$, and total energy $\rho E$ of an ideal gas. The pressure $p$ is given by the equation of state $p=(\gamma-1)(\rho E-\tfrac{1}{2}\rho|v|^{2})$, where $\gamma=1.4$ is the adiabatic constant. An invariant domain $\mathcal{G}$ of the Euler system consists of all states $u=(\rho,\rho v,\rho E)^{\top}$ such that $\rho>0$ and $p(u)>0$. The linear positivity constraint for $\rho$ can be enforced as in the scalar case. To ensure positivity preservation for $p(u)$, we adapt the limiter proposed by Abgrall et al. [abgrall-arxiv, wissocq2025] to our purposes.
@@ 5.2.1 Modified Sod’s shock tube
In the first test for the Euler equations, we consider a 1D Riemann problem corresponding to a modification [toro2009, Sec. 6.4] of Sod’s shock tube problem [sod1978]. The modified version is more challenging because a sonic point occurs within the rarefaction wave region. As a consequence, numerical methods that lack entropy stability tend to produce entropy shocks.
@@
The left boundary of the domain $\Omega=(0,1)$ is a supersonic inlet, at which the values of the primitive variables are determined by the left initial state of the Riemann problem. The boundary data $\hat{u}$ for $x=1$ may be set equal to the right initial state, since the final time $t=0.2$ is too short for the waves to reach the right boundary. The exact entropy solution features a rarefaction, a contact, and a shock wave. Its derivation can be found in [toro2009, Ch. 4].
We solve this test problem with the target scheme and its slope-limited counterpart using $N_{h}=129$ degrees of freedom per variable and $p\in\{1,2\}$. The results displayed in Figs.˜4(a) and 4(b) are in good agreement with the vanishing viscosity solution. Similarly to the test problem considered in Section˜5.1.2, the use of $\mathbb{P}_{1}$ elements and fine grids is the best option in this example. While higher-order approximations on coarser grids perform well for rarefaction waves, they are typically more diffusive at discontinuities.
@@
In Fig.˜4(c), we plot the evolution of $\eta_{\Omega}(u)=\frac{1}{1-\gamma}\int_{\Omega}\rho s(u)\mathrm{d}x$, where $s(u)=\log\left(p\rho^{-\gamma}\right)$ is the specific physical entropy. In contrast to the test case considered in Section˜5.1.2, the entropy flux $q(u)=\rho vs/(1-\gamma)$ vanishes at the boundaries, as long as the states $u(0,t)$ and $u(1,t)$ remain equal to the corresponding initial states of the Riemann problem. On the time interval for which this is the case, the total entropy $\eta_{\Omega}(u)$ of the exact vanishing viscosity solution $u$ is subject to change only due to entropy losses at the shock and/or contact. Numerically, we observe a general trend of entropy dissipation but $\eta_{\Omega}(u_{h})$ does not decrease monotonically. A probable cause of this behavior is the lack of local entropy stability, which can be cured by adding entropy viscosity or using limiter-based entropy fixes [kuzmin2020e]. However, the small wiggles in the evolution of $\eta_{\Omega}(u_{h})$ are no major concern, since no entropy shocks are observed and $\eta_{\Omega}(u_{h})\leq\eta_{\Omega}(u)$. As demonstrated already in [kuzmin2023a] in the context of scalar problems, the WENO stabilization term of our method dissipates enough entropy in most cases. Although additional entropy stabilization may still be appropriate for some scalar problems (such as the challenging but artificial KPP test in Section 5.3.2), we have never observed entropy-violating behavior of IDP finite element schemes for the Euler equations.
@@ 5.2.2 Woodward–Colella blast wave problem
The 1D blast wave problem [woodward1984] is a challenging test for the Euler equations due to strong jumps in the piecewise-constant initial pressure
@@
The initial density is $\rho_{0}\equiv 1$, and the fluid is at rest, i. e., $v_{0}\equiv 0$. The domain $\Omega=(0,1)$ has reflecting boundaries at both end points. For short time intervals, the exact solution can be obtained as for any Riemann problem of the Euler equations [toro2009, Ch. 4]. However, by the final time $t=0.038$, interactions with the boundaries and collisions of waves generated by the two Riemann problems result in the unavailability of an exact solution.
We solve this test problem numerically using $N_{h}=1001$ degrees of freedom per variable and $p\in\{1,2\}$. The density profile calculated with the target scheme is shown in Fig.˜5(a). Since no slope limiting is performed in this experiment, some negative pressures arise during the simulation run. The code does not crash only because we calculate the speed of sound using the formula $c=\sqrt{\max\{0,\gamma p/\rho\}}$. This practice is not to be recommended for non-IDP schemes, since it gives rise to consistency errors and disguises unacceptable pressure behavior. Due to round-off errors, a numerical implementation of an IDP scheme might also produce negative pressures of the order of machine precision. In this case, setting $c\coloneqq 0$ is an acceptable way to avoid complex-valued solutions. However, such situations did not occur in simulations in which we used the density limiter and the pressure limiter proposed by Abgrall et al. [abgrall-arxiv]. The slope-limited density profiles shown in Fig.˜5(b) are nonoscillatory, and no negative pressures were generated at any time.
@@ 5.2.3 Shu–Osher test / sine-shock interaction
Next, we perform simulations for the popular Shu–Osher test problem [shu1989], which uses the initial condition
@@
We test the methods under investigation using $p\in\{1,2\}$ and $N_{h}=501$ degrees of freedom per variable. Since the exact solution can be captured well by the target scheme, no violations of IDP constraints occur during the simulation run without the slope limiter. Therefore, the density and pressure fixes are not activated in practice, and the results coincide with those obtained without limiting (see Fig.˜6).
It is quite encouraging to see that the $\mathbb{P}_{2}$ version captures the post-shock oscillatory region much better than the $\mathbb{P}_{1}$ approximation on a mesh with twice as many cells. Furthermore, as is the case for the 1D blast wave problem studied in Section˜5.2.2, limiting w. r. t. global bounds does not lead to severe peak clipping effects. However, one can easily think of situations in which this issue is likely to occur, e. g., in simulations of near-vacuum states with quadratic ($\mathbb{P}_{2}$) or higher order Bernstein elements. As long as the quantities to be limited remain sufficiently far away from the global bounds of IDP constraints, optimal convergence can be achieved with high-order Bernstein elements. In the context of AFC schemes for the advection equation, numerical evidence for the validity of this claim was provided in [hajduk2020b, Sec. 7.1].
@@ 5.3 Scalar problems in 2D
To evaluate the numerical behavior of the limited CG-WENO scheme and its components in two space dimensions, we consider two additional scalar test problems, one of which is linear and the other one is nonlinear. In both cases, we restrict our studies to $\mathbb{Q}_{1}$ finite element approximations, for which the Bernstein and Lagrange basis functions coincide. Computations are performed on uniform meshes with spacing $h=\frac{1}{128}$. We use the pseudo time step $\Delta t_{e}=\Delta t_{e}^{\max}$ defined by (40) and the global time step $\Delta t=10^{-3}$, which satisfies the CFL condition (23) for both test problems. Although our choice of $\Delta t_{e}$ exceeds the threshold of Lemma 2, the high-order intermediate cell averages stay in bounds. Therefore, we deactivate the flux limiter for (44) by setting $\alpha_{ee^{\prime}}\coloneqq 1\ \forall e\in\mathcal{B}_{e}$, as proposed in Remark 1.
@@
where $\dot{u}_{h}^{L}$ denotes a low-order approximation to the time derivative $\dot{u}_{h}$. This choice is adopted because it corresponds to the second-order target employed in [kuzmin2020]. A decrease in the value of $\gamma_{e}\in[0,1]$ on the right-hand side of (58) increases the levels of low-order Rusanov dissipation (15), while decreasing the levels of high-order stabilization via the terms depending on $\Delta t_{e}$.
@@
In the WENO version, we use the linear weights $w_{l}^{e,\rm lin}=0.2$ for $l=1,\ldots,n_{e}$ and set $w_{0}^{e,\rm lin}=1-\sum_{l=1}^{n_{e}}w_{l}^{e,\rm lin}$. To study the dependence of the results on the parameter $q\geq 1$ of formula (7), we run simulations for $q\in\{3,5,10\}$. The range of approximate solutions is reported in the figures to quantify the levels of numerical diffusion and the magnitude of undershoots/overshoots.